OpenStudy (anonymous):

Fred has a savings account balance of $1,728.57. The interest rate on the account is 3.4% compounded daily. If he opened the account nine years ago, what was the value of his initial deposit?

OpenStudy (campbell_st):

are you using 365 days in the year...? and is the interest rate 3.4% per day or per year...?

OpenStudy (anonymous):

The value of his initial deposit is $14.205, because 0.034 * 9 * 365 = 121.69 and 1728.57 / 121.69 = 14.205 :)

OpenStudy (anonymous):

thank you!

OpenStudy (campbell_st):

I need to point out a small problem in the solution... compound interest formula is \[A=P(1+\frac{r}{100})^n\] using the above solution of r = 3.4% n = 9 x 365 = 3285 days and the principal being $14.20 \[A = 14.20(1+\frac{3.4}{100})^{3285}\] the future value of the investment is \[A = $3.11636 \times 10^{48}\] just a slight error...